Reduce the equation into canonical form and find the solution

xuxx +2(xy)1/2 uxy +yuyy -ux =0


1
Expert's answer
2021-01-04T18:53:50-0500

"xu_{xx}+2(xy)^{1/2}u_{xy}+yu_{yy}-u_x=0 \;\;...\pmb{(1)}\\
a=x, b=(xy)^{1/2},c=y\\
\Rightarrow b^2-ac=0\\
\Rightarrow (1) \;\; is \; parabolic.\\"

Characteristic polynomial is given by,

"\frac{dy}{dx}=\frac{b}{a}=(y/x)^{1/2}\\\;\\
\Rightarrow \frac{2}{3}y^{3/2}-\frac{2}{3}x^{3/2}=c \;\;[c\;is \; a \;constant]\\"


Define,

"\eta (x,y)=\frac{2}{3}y^{3/2}-\frac{2}{3}x^{3/2}\\
\eta_x=-\sqrt x\\
\eta_y=\sqrt y"

Choose,

"\zeta (x,y)=-x\\
\Rightarrow J=\zeta_x\eta_x-\zeta_y\eta_y=\sqrt x >0"


"w(\zeta,\eta)=u(x,y)\\
\therefore u_x=w_\zeta \zeta_y+w_\eta \eta_x=-w_\zeta-w_\eta\sqrt x\\
u_y=\sqrt y\\
u_{xx}=w_{\zeta \zeta}+\sqrt x(x+1) w_{\zeta\eta}+w_{\eta\eta}-\frac{1}{2\sqrt x}w_\eta\\
u_{xy}=-\sqrt yw_{\zeta\eta}-\sqrt{xy}w_{\eta\eta}\\
u_{yy}=yw_{\eta\eta}+\frac{1}{2\sqrt y}w_\eta"


Now from (1), the required canonical form is,

"\pmb{xw_{\zeta\zeta}+(x\sqrt x +x^2\sqrt x -2y\sqrt x )w_{\zeta\eta}+(x-y)^2w_{\eta\eta}+\frac{\sqrt y +\sqrt x}{2}w_\eta+w_\zeta=0 }"